Item 5 Item 5 The Elkmont Corporation needs to raise $63.8 million to finance its expansion...

Item 5

Item 5

The Elkmont Corporation needs to raise $63.8 million to finance its expansion into new markets. The company will sell new shares of equity via a general cash offering to raise the needed funds. The offer price is $22 per share and the company’s underwriters charge a spread of 7.5 percent. The SEC filing fee and associated administrative expenses of the offering are $1,450,000. How many shares need to be sold? (Do not round intermediate calculations and enter your answer in dollars, not millions, rounded to the nearest whole number, e.g., 1,234,567.)
Answer #1

Required proceeds net of spread=(63,800,000+SEC filing fee and associated administrative expenses)/(1-spread)

=(63,800,000+1,450,000)/(1-0.075)

=70540540.5

Hence required number of shares=70540540.5/22

=3206388 shares(Approx).